OpenStudy (anonymous):

WILL GIVE MEDAL AND COOKIE (: Which ordered pair is a solution to the equation? 2y = −5x – 4 A: (-2,3) B: (1, −4) C: (0,2) D: (−1, 1)

OpenStudy (anonymous):

is it a freshly baked cookie?

OpenStudy (anonymous):

A: 2*(3) = -5*(-2) - 4 B: 2*(-4) = -5*(1)-4 C: 2*(2) = -5*(0) -4 D: 2*(1) = -5*(1)-4 All you need to do is find out which one of them is true.

OpenStudy (anonymous):

oh sorry D: 2*(1) = -5*(-1)-4 made a typo